Please help me solve this word problem : Three years ago a woman was 3 times as old as the daughter. Their total age now is 46 years. How old was the daughter three years ago?
Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

woman = x
daughter = y
.
x-3 = 3(y-3)
x + y = 46 and therefore x = 46-y Use this value for x in the equation above:
46-y-3 = 3y-9
43-y = 3y-9
52 = 4y
y = 52/4 = 13 the daughter is 13 so 3 years ago she was 10.

Answer by ikleyn(53763)

.
It is so simple . . .

Three years ago the woman was three times as old as the daughter and the sum of their ages was 46-3*2 = 40 years.


Hence, three years ago the daughter was 10 years.     

Solved.